OpenStudy (anonymous):

h(x)+= -(x+2)(x+6) determine the vertex, x and y int and sketch a graph. Help please?

OpenStudy (eyust707):

at what points odes the function = 0?

OpenStudy (eyust707):

when x= ?? and ??

OpenStudy (anonymous):

x equals 4 i believe

OpenStudy (eyust707):

what values of x whould make h(x) 0?

OpenStudy (anonymous):

-2 and -6?

OpenStudy (eyust707):

correct

OpenStudy (eyust707):

and because of the form of this equation, we know its a parabola

OpenStudy (eyust707):

so we know its a parabola that crosses the x axis at -2 and -6

OpenStudy (eyust707):

the negitive sign indicates that its upside down

OpenStudy (eyust707):

and to find the y int set x=0

OpenStudy (eyust707):

you sketch the graph on here using the edit function.. ill check it

OpenStudy (eyust707):



OpenStudy (anonymous):

wait so which number is the vertex? -2,-6?

OpenStudy (anonymous):

h(x)+= -(x+2)(x+6) h(x)+= -(x^2+8x+12) h(x)=- -x^2-8x-12 the roots are -2,-6 vertex v=-b/2a=8/(2)(-1)=-4,plug thuis to get h(-4)=4 herefore vertex V(x,y)=V(-4,4)

OpenStudy (anonymous):